OpenStudy (anonymous):

Need help with these two vector problems! Given the following x and y components, find the magnitude and direction of the vector, counter clockwise from the x-axis: (1) x comp: -.5, y comp: 1. I found a magnitude of 1.12, but I'm having trouble with the angle. (2) x comp: -1.5 and y comp: -1; I got (1.80, 33.7 degrees, not sure if this is correct, though. I was confused with the negatives and to find the angle.

ganeshie8 (ganeshie8):

mag = \(\sqrt{x^2+y^2}\) angle = \(\tan^{-1} \frac{y}{x}\)

OpenStudy (anonymous):

I know, but I'm confused with the "counter clockwise from the x-axis" part of the problem.

ganeshie8 (ganeshie8):

angle is always measured in counter clockwise, so dont wry about it

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

btw, ur magnitudes are correct, for angle just plug it in the formula

OpenStudy (anonymous):

Is the angle for the second problem correct?

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

clearly, that 2nd vector is in 4th quadrant, so we shud get angle >180

OpenStudy (anonymous):

Hmmm, that's where I'm getting confused

ganeshie8 (ganeshie8):

yeah :) just add 180 to that angle, as tan period is \(\pi\)

ganeshie8 (ganeshie8):